OpenStudy (anonymous):

which equation has no solution? 5v+3=5v-5 2x-4=2x+7-x 2y+5-y=2y-3(y-6) 7z+9=-7z-6

OpenStudy (carolina4567):

A

OpenStudy (carolina4567):

U group like terms 5v - 5v = 3 - 5 0 = - 2 And theres no solution so there u have it .